# Provably Overwhelming Transformer Models with Designed Inputs

### v1

@@ 6.1 Experimental Setup
We train a single-layer transformer model with an embedding dimension of 512 and the BERT tokenizer (Devlin et al., 2019). The model’s architecture is outlined in section 2. We train on the AG News dataset 444see PyTorch’s documentation, using the training split with a batch size of 8, learning rate of 5e-5 using the Adam optimizer, and $20$ epochs.
We examine a few different types of input restrictions 555 The first token is always fixed to the “BOS” token which connotes the beginning of a string in the BERT tokenizer. For simplicity, we think of the BOS token as part of the fixed tokens. :
@@
  • Hamlet: We use the famous quote from Shakespeare’s Hamlet: “To be or not to be, that is the question. Whether it is a nobler in the mind to suffer the slings and arrows”
  • The Old Man and the Sea: We fix the free tokens to be a snippet from the opening chapter of Hemingway’s The Old Man and the Sea: “The blotches ran well down the sides of his face and his hands had the deep-creased scars from handling heavy fish on the cords.”
@@
In fig. 2, we show our algorithm in action for the “Hamlet” permutation class and repeated fixed string. In appendix A, we plot the worst-case deviation in fig. 3 and the peak-to-peak difference in fig. 4 for each of the input restrictions and permutation classes.
Figure 2: Worst-case deviation for the “Hamlet” permutation string and fixed repeated string. The $x$-axis is the number of total tokens and the $y$-axis is the log-scale worst-case deviation. We mark cases of provable overwhelming with a red “x.”
@@
The experimental results for worst-case deviation bounds in fig. 3 and peak-to-peak difference in fig. 4 provide interesting insights and questions. The upper-bound on worst-case deviation trend downwards as the fixed input token size is increased. But only the “repeated-string” setting is monotonically decreasing. Moreover, all plots exhibit a “phase transition:” at around 180 tokens and again around 350 tokens, the upper-bound on worst-case deviation has a sharp drop. We note that the linear program seems to be the most useful for the “AG News” dataset when paired with the The Old Man and the Sea free string. Finally, we note that in both plots, “not much” seems to happen prior to 180 tokens. Specifically, both the peak-to-peak difference and bound on worst-case deviation remain flat for the smaller fixed string.
@@ 6.2 Overwhelming Continues Through Generation
To test this we run examples of text generation where we begin with an overwhelming string $s$ and use the model to generate text. We then check if the model remains “overwhelmed”as the generated text is included as part of the fixed string. The results of these tests are plotted in fig. 5 within appendix A. One would hope the model remains overwhelmed as the fixed string is increased by the addition of the newly generated tokens. This occurs for the Hamlet free string with the repeated and random fixed strings, where as expected the worst-case deviation remains mostly constant throughout token generation. For the longer free string and the AG News fixed strings the worst-case deviation peaks after a few tokens and the model stops to be provably “overwhelmed”.
@@ Appendix A Experimental Outcomes
Figure 3: The worst-case deviation for the model for the three different input restrictions and two different permutation classes. The $x$-axis is the number of tokens and the $y$-axis is the base-10 logarithm of the worst-case deviation.
@@ Appendix B Appendix for Model Details
  • LN is the layer normalization function which for matrix $X$ performs the following column wise function for columns of $X$ (i.e. $X[:,j]$):
    $\texttt{LN}(X[i,j])=\frac{X[i,j]-\mathbb{E}[X[:,j]]}{\sqrt{\mathrm{Var}[X[:,j]]+\epsilon}}\cdot\gamma+\beta$
@@ Appendix C Proofs for Worst-Case Deviation
  • Lifting monotonicity:
    $\mathcal{W}(f;\mathcal{X})_{p}\leq\mathcal{W}(f;\mathcal{Y}\times\mathcal{Z})_{p}.$
@@
  • Lifting monotonicity: The proof follows from the definition of the lifted worst-case deviation:
    $\mathcal{W}(f;\mathcal{X})_{p}$ $=\sup_{(Y_{1},Z_{1}),(Y_{2},Z_{2})\in\mathcal{X}\subseteq\mathcal{Y}\times\mathcal{Z}}\norm{f(Y_{1},Z_{1})-f(Y_{2},Z_{2})}_{p}$
    $\leq\sup_{Y_{1},Y_{2}\in\mathcal{Y},\;Z_{1},Z_{2}\in\mathcal{Z}}\norm{f(Y_{1},Z_{1})-f(Y_{2},Z_{2})}_{p}$
    $=\mathcal{W}(f;\mathcal{Y}\times\mathcal{Z})_{p}.$
@@
Appendix D Proofs for Generic Framework
D.1 Proofs for Blowup and Shift Bounds
We first need to bound the variance and expectation of the columns of $\text{E}\cdot X$:
Lemma D.1 (Bounds on Expectation and Variance).
Let $X\in\mathbb{O}_{des}$ be an input sequence. Then for any column $j$, the expectation $\mu_{j}$ of the $j$-th column of $\text{E}\cdot X$ satisfies:
$\frac{1}{{n_{ctx}}}$ $\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n_{fix}}\text{E}[{s}_{i}]\vec{e}_{j}^{T}+n_{free}\min_{x\in[{d_{vocab}}]}(\text{E}[x]\vec{e}_{j}^{T})+\text{E}[q]\vec{e}_{j}^{T}}\right)\leq\mu_{j}$
$\leq\frac{1}{{n_{ctx}}}\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n_{fix}}\text{E}[{s}_{i}]\vec{e}_{j}^{T}+n_{free}\max_{x\in[{d_{vocab}}]}(\text{E}[x]\vec{e}_{j}^{T})+\text{E}[q]\vec{e}_{j}^{T}}\right)$
Let $\mu_{j}^{\min}$ and $\mu_{j}^{\max}$ denote the lower and upper bounds for the $j$-th column respectively. Then, we can bound the variance of the $j$-th column of $\vec{x}\text{E}$ by
$\frac{1}{{n_{ctx}}}$ $\bigg{[}\min_{\mu^{\prime}\in[\mu_{j}^{\min},\mu_{j}^{\max}]}\sum_{i=1}^{s}\mathopen{}\mathclose{{}\left(\text{E}[{s}_{i}]\vec{e}_{j}^{T}-\mu^{\prime}}\right)^{2}+n_{free}\cdot\min_{x\in[{d_{vocab}}]}\mathopen{}\mathclose{{}\left(\text{E}[x]\vec{e}_{j}^{T}-\mu^{\prime}}\right)^{2}+\mathopen{}\mathclose{{}\left(\text{E}[q]\vec{e}_{j}^{T}-\mu^{\prime}}\right)^{2}\bigg{]}$
$\leq\mathrm{Var}_{j}[(\text{E}\cdot X)[:,j]]\leq$
$\frac{1}{{n_{ctx}}}$ $\bigg{[}\sum_{i=1}^{s}\max_{\mu^{\prime}\in[\mu_{j}^{\min},\mu_{j}^{\max}]}\mathopen{}\mathclose{{}\left(\text{E}[{s}_{i}]\vec{e}_{j}^{T}-\mu^{\prime}}\right)^{2}+n_{free}\cdot\max_{x\in[{d_{vocab}}]}\mathopen{}\mathclose{{}\left(\text{E}[x]\vec{e}_{j}^{T}-\mu^{\prime}}\right)^{2}+\mathopen{}\mathclose{{}\left(\text{E}[q]\vec{e}_{j}^{T}-\mu^{\prime}}\right)^{2}\bigg{]}.$
We define $\mathrm{Var}_{j}^{\min},\mathrm{Var}_{j}^{\max}$ to be the lower and upper bounds respectively.
Proof.
The bounds on $\mu_{j}$ follow as we minimize and maximize the contribution of each free row. The proof of for variance bounds follows similarly. ∎
We now have a simple proof for lemma 4.7 restated below:
Lemma D.2 (Blowup and Shift Bounds).
We bound blowup and shift: for every $B_{j}\in\mathcal{B}_{j}$ and $S_{j}\in\mathcal{S}_{j}$
$\frac{\gamma}{\sqrt{\mathrm{Var}_{j}^{\max}+\epsilon}}\leq B_{j}\leq\frac{\gamma}{\sqrt{\mathrm{Var}_{j}^{\min}+\epsilon}}$
and
$\min\mathopen{}\mathclose{{}\left(B_{j}^{\min}\mu_{j}^{\min},B_{j}^{\min}\mu_{j}^{\max},B_{j}^{\max},\mu_{j}^{\min},B_{j}^{\max}\mu_{j}^{\max}}\right)\leq S_{j}$
$\leq\max\mathopen{}\mathclose{{}\left(B_{j}^{\min}\mu_{j}^{\min},B_{j}^{\min}\mu_{j}^{\max},B_{j}^{\max},\mu_{j}^{\min},B_{j}^{\max}\mu_{j}^{\max}}\right)$
Proof.
Given the bounds on expectation and variance, $B^{\min},B^{\max}$ follow trivially from the definition of blowup and shift sets. $S^{\min}$ (resp. $S^{\max}$) follow from a minimization (maximization) over the possible shifts given the blowup bounds. ∎
D.2 Appendix for Bounding Worst-case Deviation of $f^{enc}$
Recall the definition of blowup and shift sets from definition 4.6. Let $B^{\min},B^{\max}$ be the vectors $(B_{1}^{\min},\dots,B_{d_{emb}}^{\min})$ and $(B_{1}^{\max},\dots,B_{d_{emb}}^{\max})$ respectively. Define $S^{\min},S^{\max}$ analogously. We restate lemma 4.15 for convenience:
Lemma D.3 (Worst-case deviation of $f^{enc}$, lemma 4.15).
Let $\vec{x}\in\mathbb{O}_{des}$. Then,
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{enc};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}\leq\max_{t\in[{d_{vocab}}]}\max_{j\in[{d_{emb}}]}\max_{B\in[\vec{0},B^{\max}-B^{\min}]}\mathopen{}\mathclose{{}\left|X[t]\cdot\text{E}\cdot\mathrm{diag}(B)\cdot\vec{e}_{j}}\right|+\mathopen{}\mathclose{{}\left|S_{j}^{\max}-S_{j}^{\min}}\right|$
where the inner maximum term can be computed via a simple linear program.
Proof.
First, note that for $j\in[{d_{emb}}]$,
$\mathcal{W}(\vec{e}_{j}^{T}\cdot(f^{enc})^{T};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$ $=\max_{t,B,S,B^{\prime},S^{\prime}}\mathopen{}\mathclose{{}\left\|X[t]\cdot E\cdot\mathrm{diag}(B)+S-X[t]\cdot E\cdot\mathrm{diag}(B^{\prime})-S^{\prime}}\right\|_{\infty}$
$\leq max_{t,B,S,B^{\prime},S^{\prime}}\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left(X[t]\cdot E\cdot(\mathrm{diag}(B)-\mathrm{diag}(B^{\prime}))+S-S^{\prime}}\right)\cdot\vec{e}_{j}}\right|$
$\leq max_{t,B,B^{\prime}}\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left(X[t]\cdot E\cdot(\mathrm{diag}(B)-\mathrm{diag}(B^{\prime}))}\right)\cdot\vec{e}_{j}}\right|+|S_{j}^{\max}-S_{j}^{\min}|.$
Note then that $\mathrm{diag}(B)_{j}-\mathrm{diag}(B^{\prime})_{j}$ is constrained by $B^{\max}_{j}-B^{\min}_{j}$. We can then maximize over $j$ to get
$\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{enc};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$ $\leq\mathcal{W}(f^{enc};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{F_{\infty}}$
$\leq\max_{t,B,B^{\prime}}\max_{j}\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left(X[t]\cdot E\cdot(\mathrm{diag}(B)-\mathrm{diag}(B^{\prime}))}\right)\cdot\vec{e}_{j}}\right|+|S_{j}^{\max}-S_{j}^{\min}|$
as desired. ∎
D.3 Proofs and Subalgorithms for Attention Bounds
@@
We provide the algorithm to find the extremal values $\vec{\ell}^{\min}$ and $\vec{\ell}^{\max}$ in algorithm 4. At a high level, the algorithm computes upper and lower bounds for each position in the input sequence prior to the softmax. The bounds make use of the upper and lower bounds on the blowup and shift sets, $\mathcal{B}\mathcal{S}$.
Algorithm 4 Algorithm to find $\vec{\ell}^{\min}$ and $\vec{\ell}^{\max}$. To find the minimums and maximums over $\mathcal{B}\mathcal{S}$, we use the point-wise upper and lower bounds on the blowup and shift sets alongside the bounds obtained from algorithm 5.
@@
$\vec{\ell}^{\min}_{k}=\frac{1}{\sqrt{{d_{emb}}}}\min_{B,S\in\mathcal{B}\mathcal{S}}(\vec{e}_{q}E\cdot\mathrm{diag}(B)+S)\cdot Q\Theta_{k,{n_{ctx}}}K^{T}\cdot(\vec{e}_{{s}_{k}}E\cdot\mathrm{diag}(B)+S)^{T}$
and
$\vec{\ell}^{\max}_{k}=\frac{1}{\sqrt{{d_{emb}}}}\max_{B,S\in\mathcal{B}\mathcal{S}}(\vec{e}_{q}E\cdot\mathrm{diag}(B)+S)\cdot Q\Theta_{k,{n_{ctx}}}K^{T}\cdot(\vec{e}_{{s}_{k}}E\cdot\mathrm{diag}(B)+S)^{T}$
@@
$\vec{\ell}^{\min}_{k}=\frac{1}{\sqrt{{d_{emb}}}}\min_{t\in[{d_{vocab}}]}\min_{B,S\in\mathcal{B}\mathcal{S}}(\vec{e}_{q}E\cdot\mathrm{diag}(B)+S)\cdot Q\Theta_{k,{n_{ctx}}}K^{T}\cdot(\vec{e}_{t}E\cdot\mathrm{diag}(B)+S)^{T}$
@@
$\vec{\ell}^{\max}_{k}=\frac{1}{\sqrt{{d_{emb}}}}\max_{t\in[{d_{vocab}}]}\max_{B,S\in\mathcal{B}\mathcal{S}}(\vec{e}_{q}E\cdot\mathrm{diag}(B)+S)\cdot Q\Theta_{k,{n_{ctx}}}K^{T}\cdot(\vec{e}_{t}E\cdot\mathrm{diag}(B)+S)^{T}$
@@
$\vec{\ell}^{\min}_{n_{ctx}}=\frac{1}{\sqrt{{d_{emb}}}}\min_{B,S\in\mathcal{B}\mathcal{S}}(\vec{e}_{q}E\cdot\mathrm{diag}(B)+S)\cdot Q\Theta_{k,{n_{ctx}}}K^{T}\cdot(\vec{e}_{q}E\cdot\mathrm{diag}(B)+S)^{T}$
and
$\vec{\ell}^{\max}_{n_{ctx}}=\frac{1}{\sqrt{{d_{emb}}}}\max_{B,S\in\mathcal{B}\mathcal{S}}(\vec{e}_{q}E\cdot\mathrm{diag}(B)+S)\cdot Q\Theta_{k,{n_{ctx}}}K^{T}\cdot(\vec{e}_{q}E\cdot\mathrm{diag}(B)+S)^{T}$
@@
Algorithm 5 Simple algorithm to bound the minimum of $\vec{x}A\vec{y}^{T}$ for $X^{\min}_{i}\leq\vec{x}_{i}\leq X^{\max}_{i}$ and $Y^{\min}_{i}\leq\vec{y}_{i}\leq Y^{\max}_{i}$. To find the maximum, switch the minimum to maximum and vice versa.
  Input: $A,X^{\min},X^{\max},Y^{\min},Y^{\max}$
  $m\leftarrow 0$
  for $i=1$ to $n$ do
     for $j=1$ to $m$ do
        if $A_{ij}>0$ then
           $m\leftarrow m+A_{ij}\cdot\max\mathopen{}\mathclose{{}\left(X^{\min}_{i}Y^{\min}_{j},X^{\max}_{i}Y^{\min}_{j},X^{\min}_{i}Y^{\max}_{j},X^{\max}_{i}Y^{\max}_{j}}\right)$
        else
           $m\leftarrow m+A_{ij}\cdot\min\mathopen{}\mathclose{{}\left(X^{\min}_{i}Y^{\min}_{j},X^{\max}_{i}Y^{\min}_{j},X^{\min}_{i}Y^{\max}_{j},X^{\max}_{i}Y^{\max}_{j}}\right)$
        end if
     end for
  end for
  Return $m$
@@
In algorithm 4, for each position $k$, we compute the minimum and maximum logit for the $k$-th position by maximizing over possible blowup and shift sets and input tokens for the free tokens. Moreover, algorithm 5 computes an upper and lower bound for the restricted bilinear form which algorithm 4 uses to compute the extremal values.
The correctness of the bilinear bound in algorithm 5 follows from a series of relaxations when writing out the explicit formula for the bilinear multiplication. To prove an upper-bound, we have that
$m$ $=\sum_{i}x_{i}A_{i,j}y_{j}$
$\leq\sum_{i}\max(A_{i,j}X_{i}^{\min}Y_{i}^{\min},A_{i,j}X_{i}^{\min}Y_{i}^{\max},A_{i,j}X_{i}^{\max}Y_{i}^{\min},A_{i,j}X_{i}^{\max}Y_{i}^{\max}).$
Then, if $A_{i,j}$ is negative, we can factor the coefficient out of the maximization and flip to a minimization. Otherwise, we can simply factor $A_{i,j}$ out of the maximization. Thus, algorithm 5 gives a valid upper-bound. The lower-bound follows analogously. ∎
@@
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}$ $\leq 2\cdot({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot\max_{B,S}\norm{\text{E}[[n_{fix}]\cup\{{n_{ctx}}\},:]\cdot\mathrm{diag}(B)+S}_{{F_{\infty}}}$
$+2\cdot{\beta}_{free}^{\max}\cdot\max_{B,S}\norm{(E\cdot\mathrm{diag}(B)+S)\cdot V}_{{F_{\infty}}}+{\beta}_{free}^{\max}\norm{V}_{\infty}\cdot\max_{t\in s\cup\{q\}}\mathcal{W}(f^{enc};\{\vec{e}_{t}\}\times\mathcal{B}\mathcal{S}\}))_{\infty}$
@@
for $Y=X\text{E}\cdot\mathrm{diag}(B(X))+S(X)$. Moreover, let $\vec{p}(X)_{\text{free}}=\vec{p}(X)[(n_{fix},{n_{ctx}})]$ (i.e. the values corresponding to the free tokens) and $\vec{p}(X)_{\text{fix}}=\vec{p}(X)[[n_{fix}]\cup\{{n_{ctx}}\}]$ (the values corresponding to the fixed and query token). Now, we re-write the worst-case deviation of the attention head: for $X,X^{\prime}\in\mathbb{O}_{des}$, $(B,S),(B^{\prime},S^{\prime})\in\mathcal{B}\mathcal{S}$,
$\mathcal{W}$ $(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}\leq\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\norm{\vec{p}(X)\cdot(X\text{E}\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})\cdot(X\text{E}\cdot\mathrm{diag}(B^{\prime})+S^{\prime})V}_{\infty}$ (by definition of an attention head)
$\leq\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\bigg{\|}\vec{p}(X)_{\text{fix}}\cdot(X[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}\cdot\mathrm{diag}(B)+S)V+\vec{p}(X)_{\text{free}}\cdot(X[(n_{fix},{n_{ctx}}):,]\text{E}\cdot\mathrm{diag}(B)$
$+S)V-\vec{p}(X^{\prime})_{\text{fix}}\cdot(X^{\prime}[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}\cdot\mathrm{diag}(B^{\prime})+S^{\prime})V-\vec{p}(X^{\prime})_{\text{free}}\cdot(X^{\prime}[(n_{fix},{n_{ctx}}):,]\text{E}\cdot\mathrm{diag}(B)+S)V\bigg{\|}$
$\leq\max_{X,X^{\prime},(B,S),(B^{\prime},S^{\prime})}\bigg{\|}\vec{p}(X)_{\text{fix}}\cdot(X[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{fix}}\cdot(X^{\prime}[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}\cdot\mathrm{diag}(B^{\prime})$
$+S^{\prime})V\bigg{\|}+\bigg{\|}\vec{p}(X)_{\text{free}}\cdot(X[(n_{fix},{n_{ctx}}):,]\text{E}\cdot\mathrm{diag}(B)+S)V-\vec{p}(X^{\prime})_{\text{free}}\cdot(X^{\prime}[(n_{fix},{n_{ctx}}):,]\text{E}\cdot\mathrm{diag}(B)+S)V\bigg{\|}.$ (by triangle inequality)

### v2

@@ 6.1 Experimental Setup
We train a single-layer transformer model with causal masking, embedding dimension of 512, and the BERT tokenizer (Devlin et al., 2019). The model’s architecture is outlined in section 2. We train on the AG News dataset 333see PyTorch’s documentation, using the training split with a batch size of 8, learning rate of 5e-5 using the Adam optimizer, and $20$ epochs.
We examine a few different types of input restrictions 444 The first token is always fixed to the “BOS” token which connotes the beginning of a string in the BERT tokenizer. For simplicity, we think of the BOS token as part of the fixed tokens. :
@@
  • The Old Man and the Sea: We fix the free tokens to be a snippet from the opening chapter of Hemingway’s The Old Man and the Sea: “The blotches ran well down the sides of his face”
  • Hamlet: We use the famous quote from Shakespeare’s Hamlet: “To be, or not to be, that is the question: Whether ’tis nobler in the mind to suffer The slings and arrows of outrageous fortune,”
@@
In appendix A, we plot the worst-case deviation in fig. 2 and the peak-to-peak difference in fig. 3 for each of the input restrictions and permutation classes.
@@
The experimental results for worst-case deviation bounds in fig. 2 and peak-to-peak difference in fig. 3 provide interesting insights and questions. The upper-bound on worst-case deviation trend downwards as the fixed input token size is increased. But only the “repeated-string” setting is monotonically decreasing. We note that the linear program seems to give the most advantage over the naive approach for the “Hamlet” free string. This makes sense, as the “Hamlet” string is much longer than the “Old Man and the Sea” string, and thus the linear program has more room to give a better bound.
@@ 6.2 Overwhelming Continues Through Generation
To test this we run examples of text generation where we begin with an overwhelming string $s$ and use the model to generate text. We then check if the model remains “overwhelmed”as the generated text is included as part of the fixed string. The results of these tests are plotted in fig. 4 within appendix A. One would hope the model remains overwhelmed as the fixed string is increased by the addition of the newly generated tokens. This seems to occur for both strings: once overwhelming provably occurs, the model remains “overwhelmed”for the remainder of the generated tokens.
@@ Appendix A Experimental Outcomes
Figure 2: The worst-case deviation for the model for the three different input restrictions and two different permutation classes. The $x$-axis is the number of tokens and the $y$-axis is the upper-bound on worst-case deviation.
@@ Appendix B Appendix for Model Details
  • LN is the layer normalization function which for matrix $X$ performs the following row-wise (token-wise) function for rows of $X$ (i.e. $X[i,:]$):
    $\texttt{LN}(X[i,j])=\frac{X[i,j]-\mathbb{E}[X[i,:]]}{\sqrt{\mathrm{Var}[X[i,:]]+\epsilon}}\cdot\gamma+\beta$
@@
    Further, we will restrict our attention to causal attention, which means that the attention matrix is upper triangular. Formally,
    $\texttt{AttnH}(X)=\text{softmax}\mathopen{}\mathclose{{}\left(\frac{\operatorname{RoPE}(XQ)\cdot\operatorname{RoPE}(K^{T}X^{T})}{\sqrt{{d_{emb}}/H}}+M}\right)\cdot XV$
    where
    $M_{ij}=\begin{cases}0&i\leq j\\ -\infty&i>j.\end{cases}$
@@ Appendix C Proofs for Worst-Case Deviation
Appendix D Proofs for section 4
D.1 Proofs and Subalgorithms for Attention Bounds
@@
We provide the algorithm to find the extremal values $\vec{\ell}^{\min}$ and $\vec{\ell}^{\max}$ in algorithm 4. At a high level, the algorithm computes upper and lower bounds for each position in the input sequence prior to the softmax.
Algorithm 4 Algorithm to find $\vec{\ell}^{\min}$ and $\vec{\ell}^{\max}$.
@@
$\vec{\ell}^{\min}_{k}=\vec{\ell}^{\max}_{k}=\frac{1}{\sqrt{{d_{emb}}}}\vec{e}_{q}\text{E}_{\texttt{LN}}\cdot Q\cdot\Theta_{k,{n_{ctx}}}K^{T}\cdot\text{E}_{\texttt{LN}}^{T}\cdot\vec{e}_{{s}_{k}}^{T}$
@@
$\vec{\ell}^{\min}_{k}=\frac{1}{\sqrt{{d_{emb}}}}\min_{t\in[{d_{vocab}}]}\vec{e}_{q}\cdot\text{E}_{\texttt{LN}}\cdot Q\Theta_{k,{n_{ctx}}}\cdot K^{T}\cdot(\vec{e}_{t}\cdot\text{E}_{\texttt{LN}})^{T}$
@@
$\vec{\ell}^{\min}_{k}=\frac{1}{\sqrt{{d_{emb}}}}\max_{t\in[{d_{vocab}}]}\vec{e}_{q}\cdot\text{E}_{\texttt{LN}}\cdot Q\Theta_{k,{n_{ctx}}}\cdot K^{T}\cdot(\vec{e}_{t}\cdot\text{E}_{\texttt{LN}})^{T}$
@@
$\vec{\ell}^{\min}_{n_{ctx}}=\vec{\ell}^{\max}_{n_{ctx}}=\frac{1}{\sqrt{{d_{emb}}}}(\vec{e}_{q}\cdot\text{E}_{\texttt{LN}})\cdot Q\cdot\Theta_{k,{n_{ctx}}}K^{T}\cdot(\vec{e}_{q}\cdot\text{E}_{\texttt{LN}})^{T}$
@@
In algorithm 4, for each position $k$, we compute the minimum and maximum logit for the $k$-th position by maximizing over possible input tokens for the free tokens. Note that the fixed and query tokens have constant logit value, and so the minimum and maximum logit values equal each other. ∎
@@
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des})_{\infty}$ $\leq({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot\norm{\text{E}_{\texttt{LN}}[[n_{fix}]\cup\{{n_{ctx}}\},:]\cdot V\cdot\mathrm{Unembed}}_{{F_{\infty}}}+2\cdot{\beta}_{free}^{\max}\cdot\norm{\text{E}_{\texttt{LN}}\cdot V\cdot\mathrm{Unembed}}_{{F_{\infty}}}.$
@@
For notational simplicity, we will absorb the $\mathrm{Unembed}$ matrix into the value matrix $V$. So, here we will write $V=V^{\prime}\cdot\mathrm{Unembed}$ where $V^{\prime}$ is the value matrix in the attention head.
@@
for $Y=X\text{E}_{\texttt{LN}}$. Moreover, let $\vec{p}(X)_{\text{free}}=\vec{p}(X)[(n_{fix},{n_{ctx}})]$ (i.e. the values corresponding to the free tokens) and $\vec{p}(X)_{\text{fix}}=\vec{p}(X)[[n_{fix}]\cup\{{n_{ctx}}\}]$ (the values corresponding to the fixed and query token). Now, we re-write the worst-case deviation of the attention head: for $X,X^{\prime}\in\mathbb{O}_{des}$,
$\mathcal{W}$ $(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}\leq\max_{X,X^{\prime}}\norm{\vec{p}(X)\cdot(X\text{E}_{\texttt{LN}}\cdot V)-\vec{p}(X^{\prime})\cdot(X^{\prime}\cdot V}_{\infty}$ (by definition of an attention head)
$\leq\max_{X,X^{\prime}}\bigg{\|}\vec{p}(X)_{\text{fix}}\cdot(X[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}_{\texttt{LN}})V+\vec{p}(X)_{\text{free}}\cdot(X[(n_{fix},{n_{ctx}}):,]\text{E}_{\texttt{LN}})V$
$-\vec{p}(X^{\prime})_{\text{fix}}\cdot(X^{\prime}[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}_{\texttt{LN}})V-\vec{p}(X^{\prime})_{\text{free}}\cdot(X^{\prime}[(n_{fix},{n_{ctx}}):,]\text{E}_{\texttt{LN}})V\bigg{\|}$
$\leq\max_{X,X^{\prime}}\bigg{\|}\vec{p}(X)_{\text{fix}}\cdot(X[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}_{\texttt{LN}})V-\vec{p}(X^{\prime})_{\text{fix}}\cdot(X^{\prime}[[n_{fix}]\cup\{{n_{ctx}}\}:,]\text{E}_{\texttt{LN}}\cdot)V\bigg{\|}$
$+\bigg{\|}\vec{p}(X)_{\text{free}}\cdot(X[(n_{fix},{n_{ctx}}):,]\text{E}_{\texttt{LN}})V-\vec{p}(X^{\prime})_{\text{free}}\cdot(X^{\prime}[(n_{fix},{n_{ctx}}):,]\text{E}_{\texttt{LN}})V\bigg{\|}.$ (by triangle inequality)